# GraphEx: A Graph-based Extraction Method for Advertiser Keyphrase Recommendation

### v3

@@ I-B Scope and Contributions
In this work, we focus on the challenges mentioned in the previous section. We limit ourselves to retrieving keyphrases based on item’s title and belonging to the items’ categorical populace under budgetary constraints, especially those keyphrases that are actively and frequently searched by buyers. The extraction is done in an unsupervised setting where the keyphrases for each item are unknown during training. In fact, we restrict the curation of keyphrases (more details in III-B) to include only those that have a high search volume (number of searches made by buyers) based on buyer searches. XMC models suffer from the biases mentioned in Section I-A2 and as a result tend to recommend tail keyphrases (Section I-A1). Our distinction is that by using categorical population dynamics of keyphrases, we decouple the keyphrases from item click-based engagement. This allows us to retain the essential bias towards head keyphrases (attractive to advertisers) while getting rid of the negative bias (against non-popular items which are the main target of advertisement).
@@ III-B Dataset Curation
We aggregate our keyphrase datasets and their category associations from the search logs generated during buyer sessions on eBay.com. The keyphrases that buyers input during the search sessions are curated based on certain criteria, which we discuss here. The categorization of items are in form of a tree. The root level category is called as meta category (top level category) which branches down to the Leaf Category. eBay’s search engine Cassini shows a sufficient number of items (Recall Count) for each input query in its search results. Cassini determines the leaf category of the keyphrase and it is the same as the top-ranked item’s leaf category (lowest-level product categorization). Note that the item-keyphrase click-based associations are not curated into our datasets.
We restrict the number of curated keyphrases by only considering those that are heavily searched by the buyers. The number of times a keyphrase is queried is termed as (Search Count). The search count is used to explicitly control the proportions of head keyphrases in our dataset. The absolute values of Recall Count and Search Count are not essential in fact, an ranking (integer value) also works. All the unique keyphrases are aggregated for each meta category and are grouped for each leaf category within the meta category. Each keyphrase is associated with a Search Count and a Recall Count. Note that a keyphrase can be duplicated across different Leaf Categories.
@@ IV-A Setup and Datasets
GraphEx is implemented for multi-core systems without requiring any GPUs. Its inference part is implemented on C++ ($\geq$ g++-9.3.0) using OpenMP threading with Python wrappers using pybind11. The construction part is implemented in Python ($\geq$ 3.7); due to its lightweight approach since the construction does not require large resources and takes much less time. We used a system with 4 Intel Xeon Gold 6230R CPUs with 2 sockets each containing 20@2.10GHz cores, and 500 GB of RAM for the analysis. GraphEx employs coarse-grained multithreading, assigning each input’s inference to an individual thread. We launched 20 threads with compact pinning to occupy only a single socket sufficient for our dataset size.
@@ IV-B Experimentation Details
The data is collected from search logs for the duration of one year for XMC models and 6 months for GraphEx. For XMC, the item-keyphrase pairs are constrained based on their co-occurrence count, number of buyer clicks/purchases, etc. The curated unique keyphrase count shown in the third column of table II contains both the head and tail keyphrases and is incorporated by XMC models. On the other hand, GraphEx’s data curation for training, aggregates keyphrases without looking at any (clicked-based) association with the items. It restricts the keyphrases [^5] Shown in the right most column of table II. to contain a higher number of head and a lower number of tail keyphrases using the curation process described in Section III-B. Generally, keyphrases that on an average weren’t searched at least once per day were filtered for GraphEx [^6] This restriction was relaxed for CAT_3 that didn’t have sufficient keyphrases..
@@ IV-C Evaluation Framework
We described here our evaluation framework that tackles the challenges in Section I-A3. Summarizing the two major problems are 1) Lack of ground truths due to sparse data and MNAR-led biases in click data and 2) Model Convergence due to similar training data and evaluation. Ideally, metrics should compare the relevancy of the predictions to the input text without limiting the comparison to a set of predefined labels/keyphrases. However, it is difficult to determine the relevance of predictions without any prior labels or the absence of negative labels [^7] Absence of click does not imply irrelevance, see Section I-A2.
@@
Once a set of keyphrases is determined as relevant for an item, we filter the keyphrases through the high Search Count threshold. This threshold is determined as the 90th percentile of the descending order of search counts of all unique keyphrases in the category such that 10% of the unique keyphrases are above this limit. The keyphrases whose Search Counts are above the threshold are considered as Relevant Head Keyphrases [^10] High Search Count. otherwise they are considered as Relevant Tail Keyphrases. [^11] This evaluation framework is only for offline analysis and none of the model’s recommendations to the seller are filtered in this way as it would be infeasible to do so due to the latency of LLMs. One intuition behind this evaluation is that head keyphrase determined as irrelevant by the AI will be ineffective — even though buyers search the keyphrases in large volume, they mostly will not click on the corresponding item.
@@
Due to the varying number of predictions by each model, we use one set of metrics to compare the relevant and head keyphrases within each model (RP and HP) and another set to compare between different models (RRR and RHR).
@@ IV-D2 Diversity Comparison
Figure 5: Venn diagram representing the different recall sources and their keyphrases. The intersecting regions depect the keyphrases from multiple recall sources, while the non-intersecting crossed out regions depict the keyphrases belonging solely to that recall source — which contribute to the incremental impact which each recall source brings.
Therefore, the final metric to compare with other models is the exclusive diversity of GraphEx’s predictions. For this evaluation, as illustrated in Figure 5, we first separate out the unique head keyphrases recommended by each model that are relevant to the item. Table IV shows the relative amount of relevant exclusive head keyphrases of each model to the relevant exclusive head keyphrases of GraphEx (averaged per item). It is evident that GraphEx recommends the highest amount of diverse relevant head keyphrases in contrast to any other model.
@@ IV-F1 Alignment functions
TABLE VI: Table showing the relevant proportion (RP) of Word Match Ratio (WMR), Jaccard Coefficient(JAC) and Label Title Alignment (LTA) when used in GraphEx algorithm for the 3 categories.
@@ IV-G Execution Results
It is important for the models to attain the real-time recommendation and model refresh goals as described in section I-A5. We compare only the XMC models with GraphEx as the REs and SLs (except SL-emb) are simple retrieval techniques implemented in the Spark/Hadoop ecosystem while model inferencing is more complex technique. We examine the models based on Inference Latencies, Model Sizes, and Training times.  [^12] SL-emb inference stages are complex, embedding generation occurs in GPU whereas ANN is done in CPU, thus it is difficult to compare the inference latencies with other models.
@@
For near real-time recommendation, the Inference Latency of a single input should be in milliseconds. The Figure 6(a) compares the per-input inference latency of the XMC models and GraphEx. The latencies for each model are computed by amortizing the time taken for prediction over the entire test set. We can see that all the models are within the required limit of 10 ms, but fastText takes more time for a prediction. Graphite and GraphEx’s latencies are comparable for the smaller categories (CAT_2 and CAT_3). The performance of GraphEx is superior, attaining up to 17×17$\times$17 × and up to 13×13$\times$13 × more speed up in contrast to fastText and Graphite on CAT_1. If we infer 20 million items in CAT_1, GraphEx will result in energy savings of 11 hours and 8.5 hours with respect to fastText and Graphite, respectively.
The Figure 6(b) compares the storage sizes of the above models. The fastText model requires significantly more storage across all categories because of the extensive weight matrix and the word embeddings it maintains. This is the case even after reducing the model size during training to enhance precision for production. Graphite occupies substantial space for the large category CAT_1 but has a comparable size to GraphEx for other categories. GraphEx occupies the least size for its models even after constructing graphs for multiple leaf categories. The training times of fastText run into $>4$ hours for all categories with bigger categories running for days and include multiple epochs and autotuning phases. Graphite has a graph-based construction step that takes around $1-6$ minutes while GraphEx takes $<1$ minutes on all the categories. This is due to the curtailment of the training data to head keyphrases and implementation of the construction step that efficiently constructs and stores the model.
@@ IV-H Production Engineering Architecture
Figure 7: GraphEx Batch/NRT Serving Architecture
GraphEx’s batch inference is done using eBay’s machine learning platform Krylov[40] and runs on a single node with 70 cores and 900 GB RAM. GraphEx inference is so fast that the time required to run on a space of 200 million items is just 1.5 hrs. This is a huge improvement over fastText and Graphite which take 1.75 days and 1.5 days respectively. Another batch job in Spark joins these sources in Hadoop and injects them into a Key-Value store (NuKV), which is then called by the eBay platform’s inference api and served to eBay’s sellers. This architecture can scale to billions of items and hundreds of billions of keyphrases that serve eBay’s platform. The architecture is illustrated in Figure  7. Due to the nature of its algorithm, the GraphEx model is bounded by the label space on which it trains. However, since GraphEx training is as inexpensive as Graphite, the model can train in a matter of minutes even for very large categories, making it ideal for daily model refresh. This makes it possible for GraphEx to cater to newer keyphrases that arise every day. This is a huge improvement over fastText which takes a day or more to train on these large categories and has a monthly refresh schedule.

### v4

@@ I-B Scope and Contributions
In this work, we focus on the challenges mentioned in the previous section. We limit ourselves to retrieving keyphrases based on item’s title and belonging to the items’ categorical populace under budgetary constraints, especially those keyphrases that are actively and frequently searched by buyers. The extraction is done in an unsupervised setting where the keyphrases for each item are unknown during training. In fact, we restrict the curation of keyphrases (more details in III-B) to include only those that have a high search volume (number of searches made by buyers) based on buyer searches. XMC models suffer from the biases mentioned in Section I-A2 and as a result tend to recommend tail keyphrases (Section I-A1). Our distinction is that by using categorical population dynamics of keyphrases, we decouple the keyphrases from item click-based engagement. This allows us to retain the essential bias towards head keyphrases (attractive to advertisers) while getting rid of the negative bias (against non-popular items which are the main target of advertisement) which helps us solve Challenge I-A2.
@@ III-B Dataset Curation
We aggregate our keyphrase datasets and their category associations from the search logs generated during buyer sessions on eBay.com. The keyphrases that buyers input during the search sessions are curated based on certain criteria, which we discuss here. The categorization of items are in form of a tree. The root level category is called as meta category (top level category) which branches down to the Leaf Category. eBay’s search engine Cassini shows a sufficient number of items (Recall Count) for each input query in its search results. Cassini determines the leaf category of the keyphrase and it is the same as the top-ranked item’s leaf category (lowest-level product categorization). Note that, as mentioned in Section I-B the item-keyphrase click-based associations are not curated into our datasets, thus avoiding related biases and addressing Challenge I-A2.
We restrict the number of curated keyphrases by only considering those that are heavily searched by the buyers (head keyphrases). This helps address Challenge I-A1 by enabling GraphEx’s final recommendations to have higher realizations adhering to budget constraints. The number of times a keyphrase is queried is termed as (Search Count). The search count is used to explicitly control the proportions of head keyphrases in our dataset. The absolute values of Recall Count and Search Count are not essential in fact, an ranking (integer value) also works. All the unique keyphrases are aggregated for each meta category and are grouped for each leaf category within the meta category. Each keyphrase is associated with a Search Count and a Recall Count. Note that a keyphrase can be duplicated across different Leaf Categories. The curated dataset is used to construct GraphEx’s graph (described later) from keyphrases that buyers actively search tackling Challenge I-A4.
@@ III-E1 Enumeration Step
The Enumeration step aids in resolving the permutation problem laid out in Section III-A. Along with the curation process it retrieves keyphrases that are active and impactful, focusing on our targeting goal, as explained in Challenge I-A4.
@@ III-E2 Ranking Step
This results in more relevant keyphrases appearing at the top of recommendations, effectively addressing Challenge I-A1, especially when there is a limitation on the number of recommendations.
@@ III-F Implementation Details
GraphEx’s inference algorithm and implementation have been designed for efficiency, aligning with our objectives for Challenge I-A5. In this section, we examine the time complexity of the inference algorithm and explain how certain steps in its implementation are optimized to achieve real-time latency.
@@ IV-A Setup and Datasets
GraphEx is implemented for multi-core systems without requiring any GPUs. Its inference part is implemented on C++ ($\geq$ g++-9.3.0) using OpenMP threading with Python wrappers using pybind11. The construction part is implemented in Python ($\geq$ 3.7); due to its lightweight approach since the construction does not require large resources and takes much less time. We used a system with 4 Intel Xeon Gold 6230R CPUs with 2 sockets each containing 20@2.10GHz cores, and 500 GB of RAM for the analysis. A single machine is utilized here for test standardization. GraphEx employs coarse-grained multithreading, assigning each input’s inference to an individual thread. We launched 20 threads with compact pinning to occupy only a single socket sufficient for our dataset size.
@@ IV-B Experimentation Details
The data is collected from search logs for the duration of one year for XMC models and 6 months for GraphEx. For XMC, the item-keyphrase pairs are constrained based on their co-occurrence count, number of buyer clicks/purchases, etc. The curated unique keyphrase count shown in the third column of table II contains both the head and tail keyphrases and is incorporated by XMC models. On the other hand, GraphEx’s data curation for training, aggregates keyphrases without looking at any (clicked-based) association with the items. It restricts the keyphrases to contain a higher number of head and a lower number of tail keyphrases using the curation process described in Section III-B. Generally, keyphrases that on an average weren’t searched at least once per day were filtered for GraphEx [^5] The constraint was eased for CAT_3 due to a lack of enough keyphrases..
@@ IV-C Evaluation Framework
We describe here our evaluation framework that tackles the challenges in Section I-A3. Summarizing the two major problems are 1) Lack of ground truths due to sparse data and MNAR-led biases in click data and 2) Model Convergence due to similar training data and evaluation. Ideally, metrics should compare the relevancy of the predictions to the input text without limiting the comparison to a set of predefined labels/keyphrases. However, it is difficult to determine the relevance of predictions without any prior labels or the absence of negative labels. Thus, directly using traditional metrics like precision and recall which require ground truths incorporates the biases described in Section I-A3.
@@
Keyphrases identified as pertinent to an item undergo filtering through a high Search Count threshold, which is set at the 90th percentile of search counts for all unique keyphrases in the category, ensuring 10% exceed this limit. Those surpassing this threshold are labeled as Relevant Head Keyphrases, whereas others are deemed Relevant Tail Keyphrases. [^8] This evaluation method is solely for offline analysis; the model’s seller recommendations are not filtered this way due to LLM latency constraints. The rationale is that keyphrases deemed irrelevant by AI will not perform effectively; buyers may frequently search them, but seldom click on the corresponding item.
@@
Due to the varying number of predictions by each model, we use one set of metrics to compare the relevant and head keyphrases within each model (RP and HP) and another set to compare between different models (RRR and RHR). By disassociating from traditional metrics we avoid relying on ground truth biases thus addressing Challenge I-A3.
@@ IV-D2 Diversity Comparison
Figure 5: Venn diagram illustrating various recall sources and their associated keyphrases. The intersecting areas show keyphrases shared by multiple sources, whereas the crossed-out, non-overlapping sections represent keyphrases unique to each source, highlighting their distinct contributions.
Therefore, the final metric to compare with other models is the exclusive diversity of GraphEx’s predictions. For this evaluation, as illustrated in Figure 5, we first separate out the unique head keyphrases recommended by each model that are relevant to the item. Table IV shows the relative amount of relevant exclusive head keyphrases of each model to the relevant exclusive head keyphrases of GraphEx (averaged per item). It is evident that GraphEx recommends the highest amount of diverse relevant head keyphrases in contrast to any other model. Since this evaluation does not depend on ground truths, we can better assess the number of unique keyphrases that each model recommends, thus also assessing model convergence, which is a key challenge in I-A3.
@@ IV-F1 Alignment functions
TABLE VI: Relevant proportion (RP) of Word Match Ratio (WMR), Jaccard Coefficient (JAC) and Label Title Alignment (LTA) when used in GraphEx algorithm for the 3 categories.
@@ IV-G Execution Results
It is important for the models to attain the real-time recommendation and model refresh goals as described in section I-A5. We compare only the XMC models with GraphEx as the REs and SLs (except SL-emb) are simple retrieval techniques implemented in the Spark/Hadoop ecosystem while model inferencing is more complex technique. We examine the models based on Inference Latencies, Model Sizes, and Training times.  [^9] The inference phases of SL-emb are intricate, as embedding creation is processed on the GPU while ANN operations occur on the CPU, complicating latency comparisons with alternative models.
@@
For near real-time recommendation, the Inference Latency of a single input should be in milliseconds. Figure 6(a) compares the per input inference latency of the XMC models and GraphEx. The latencies for each model are computed by amortizing the time taken for prediction over the entire test set. We can see that all the models are within the required limit of 10 ms, but fastText takes more time for a prediction. Graphite and GraphEx’s latencies are comparable for the smaller categories (CAT_2 and CAT_3). The performance of GraphEx is superior, attaining up to 17×17$\times$17 × and up to 13×13$\times$13 × more speed up in contrast to fastText and Graphite on CAT_1. If we infer 20 million items in CAT_1, GraphEx will result in energy savings of 11 hours and 8.5 hours with respect to fastText and Graphite, respectively.
Figure 6(b) evaluates model storage demands. fastText consumes considerably more memory due to its weight matrix and word embeddings, despite a size reduction later to boost production accuracy. Graphite’s storage demand is sizeable for CAT_1 but aligns with GraphEx in other cases. GraphEx maintains minimal storage requirements, even with multiple leaf category graphs. FastText training surpasses $4$ hours, extending to several days for larger groups, encompassing extensive epochs and autotuning. Graphite’s graph construction spans $1-6$ minutes, whereas GraphEx completes in under $1$ minute, aided by focused data training and effective model construction execution.
Figure 7: GraphEx Batch/NRT Serving Architecture
@@ IV-H Production Engineering Architecture
GraphEx’s batch inference leverages eBay’s Krylov machine learning platform[40] and is executed on a single node featuring 70 cores and 900 GB RAM. The inference task for 200 million items for one site completes in merely 1.5 hours, vastly surpassing the performance of fastText and Graphite, which require 1.75 and 1.5 days, respectively. [^10] In production, GraphEx can operate in a distributed manner, adapting to resource availability and workload. It achieves parallelization by replicating the constructed graph across all machines and evenly distributing the inference load for multiple sites. An additional Spark batch process utilizes Hadoop to integrate these sources into NuKV, a Key-Value store accessed via eBay’s inference API, subsequently serving sellers on the platform. This system, depicted in Figure  7, scales efficiently to handle billions of items and keyphrases. GraphEx’s algorithm is limited by the label space it uses for training but remains cost-effective like Graphite, requiring only minutes to train even large category models, enabling daily model refresh. This allows adaptation to new keyphrases arising daily, significantly improving upon fastText, which demands over a day for training large categories and updates monthly.